# TELL-TALE: Task Efficient LLMs with Task Aware Layer Elimination

### v1

@@ 5 Results
Dataset LLaMA 3.1 8B (zero-shot) Qwen 2.5 7B (zero-shot)
Baseline Best Model BSBA Baseline Best Model BSBA
Perf. Perf. #D Sp. Perf. #D Sp. Perf. Perf. #D Sp. Perf. #D Sp.
ARC-Easy 87.00 90.55(+4.08% $\uparrow$) 5 1.3 87.82 8 1.4 90.04 94.40(+4.84% $\uparrow$) 3 1.2 90.08 7 1.5
ARC-Challenge 75.86 78.62(+3.63% $\uparrow$) 4 1.3 76.90 7 1.4 86.55 92.00(+6.45% $\uparrow$) 2 1.3 86.55 6 1.4
BoolQ 85.00 86.20(+1.40% $\uparrow$) 3 1.1 85.70 7 1.4 81.90 83.90(+2.44% $\uparrow$) 4 1.3 82.70 5 1.2
MMLU 54.87 59.90(+9.17% $\uparrow$) 1 1.1 54.87 9 1.4 68.10 71.00(+4.26% $\uparrow$) 5 1.2 68.13 6 1.3
CommonQA 72.20 75.30(+4.29% $\uparrow$) 3 1.2 73.10 6 1.3 80.30 84.40(+5.11% $\uparrow$) 2 1.1 80.50 6 1.3
Winogrande 53.83 56.67(+5.28% $\uparrow$) 4 1.3 53.83 12 1.5 62.04 67.25(+8.40% $\uparrow$) 3 1.2 62.19 6 1.5
BIG-Bench 75.20 83.60(+11.17% $\uparrow$) 5 1.3 75.20 11 1.6 79.20 81.60(+3.03% $\uparrow$) 6 1.5 81.60 6 1.5
GSM8K-HARD 15.07 37.08(+146.05% $\uparrow$) 1 1.1 17.31 6 1.4 7.8 26.8(+243.58% $\uparrow$) 2 1.1 8.99 5 1.2
Math500 20.50 26.00(+26.83% $\uparrow$) 1 1.1 26.00 3 1.2 18.00 27.00(+50.0% $\uparrow$) 2 1.1 21.00 4 1.2
Dataset Lucie 7B (zero-shot) Mistral 7B (zero-shot)
Baseline Best Model BSBA Baseline Best Model BSBA
Perf. Perf. #D Sp. Perf. #D Sp. Perf. Perf. #D Sp. Perf. #D Sp.
ARC-Easy 72.45 76.55(+5.66% $\uparrow$) 6 1.3 72.55 13 1.8 81.02 83.45(+4.23% $\uparrow$) 5 1.3 81.09 9 1.5
ARC-Challenge 48.00 53.79(+12.06% $\uparrow$) 7 1.4 51.38 11 1.8 72.20 74.83(+3.64% $\uparrow$) 6 1.4 72.41 8 1.4
BoolQ 53.70 77.50(+44.32% $\uparrow$) 5 1.3 60.60 19 2.8 80.36 83.20(+3.53% $\uparrow$) 6 1.2 80.60 10 1.5
MMLU 21.36 42.98(+101.2% $\uparrow$) 8 1.4 22.73 22 3.3 52.73 57.81(+9.63% $\uparrow$) 2 1.0 52.91 8 1.4
CommonQA 55.50 69.70(+25.59% $\uparrow$) 3 1.5 57.10 17 2.6 57.32 61.40(+7.12% $\uparrow$) 4 1.2 57.40 7 1.3
Winogrande 54.20 57.80(+6.64% $\uparrow$) 5 1.2 54.30 15 1.9 52.55 58.80(+11.53% $\uparrow$) 10 1.6 53.43 13 1.8
BIG-Bench 69.60 77.20(+9.84% $\uparrow$) 9 1.2 72.00 15 1.6 70.00 76.40(+9.14% $\uparrow$) 9 1.3 72.80 11 1.4
GSM8K-HARD 14.20 17.80(+25.35% $\uparrow$) 1 1.1 17.40 3 1.1 11.24 19.10(+69.92% $\uparrow$) 2 1.1 15.73 4 1.2
Math500 19.00 27.00(+42.11% $\uparrow$) 2 1.1 26.00 3 1.2 8.00 16.00(+100% $\uparrow$) 1 1.0 10.00 4 1.2
Dataset Qwen 2.5 0.5B (zero-shot)
Baseline Best Model BSBA
Perf. Perf. #D Sp. Perf. #D Sp.
ARC-Easy 40.00 60.91(+48.49% $\uparrow$) 3 1.1 48.36 5 1.4
ARC-Challenge 35.52 40.34(+13.57% $\uparrow$) 1 1.1 37.24 4 1.5
BoolQ 62.30 67.20(+7.87% $\uparrow$) 5 1.4 66.20 6 1.5
MMLU 31.48 39.97(+26.96% $\uparrow$) 2 1.1 33.90 5 1.4
CommonQA 42.40 49.10(+15.80% $\uparrow$) 2 1.3 44.00 3 1.4
Winogrande 49.86 51.88(+4.51% $\uparrow$) 5 1.3 49.87 17 3.9
BIG-Bench 72.40 73.60(+1.66% $\uparrow$) 2 1.2 73.60 2 1.2
GSM8K-HARD 6.74 11.24(+66.77% $\uparrow$) 1 1.2 8.99 2 1.2
Math500 8.00 12.00(+50% $\uparrow$) 1 1.1 9 2 1.1
Table 1: Performance comparison across language models under 0-shot evaluation. We report accuracy (Perf.), number of dropped layers (#D), and relative inference speedup (Sp.). Percentage gain = $\frac{\text{Acc}_{\text{Best}}-\text{Acc}_{\text{Baseline}}}{\text{Acc}_{\text{Baseline}}}\times 100$. Best accuracy is highlighted in bold; BSBA shows balanced trade-offs.
Iterative pruning trajectories. Figure 2 visualizes the iterative layer-pruning process for LLaMA 3.1 8B. Each curve tracks accuracy as layers are progressively removed. The $\star$ denotes the best-performing pruned model (Best), while the $\square$ highlights the Best Speedup with Baseline Accuracy (BSBA) model—the pruned configuration achieving maximum inference speedup without falling below baseline accuracy.
From these trajectories, three consistent patterns emerge: (i) Tale identifies compressed models that outperform the original across diverse tasks, with $\star$ markers lying strictly above baseline. (ii) Accuracy improvements persist across multiple pruning steps before diminishing returns, showing that substantial redundancy exists even in carefully tuned pretrained models. (iii) Pruning dynamics are task-specific: datasets such as ARC-Easy and MMLU tolerate deeper pruning while continuing to improve, whereas reasoning-heavy tasks like GSM8K-Hard converge earlier, reflecting heterogeneous layer importance across domains.
Best vs. BSBA models. Table 1 compares baseline models against their pruned counterparts under both Best and BSBA configurations. Across all benchmarks, the Best models yield consistent accuracy gains—up to $146\%$ (LLaMA 8B on GSM8K-Hard), $101\%$ (Lucie 7b on MMLU) and $244\%$ (Qwen 7B on GSM8k-Hard)—while also delivering moderate speedups. BSBA models, by construction, trade smaller gains in accuracy for more aggressive speedups, offering practical operating points when inference cost is the dominant concern.
Few-shot setting. We tested Tale under the few-shot regime for Lucie and LLaMA models (Appendix Tables 45). Few-shot prompting improves baselines on reasoning tasks such as GSM8K and Math500, yet Tale-pruned variants still achieve higher accuracy in nearly all settings. This shows that pruning-induced improvements are largely complementary to gains from in-context learning.
Takeaways. Tale consistently uncovers Pareto-optimal models along the accuracy–efficiency frontier. By balancing task fidelity with computational savings, it enables both accuracy-focused and efficiency-focused deployment. The observed diversity in pruning profiles across datasets underscores the importance of adaptive pruning, rather than one-size-fits-all heuristics, for effective model compression. A tunable selection metric for choosing among candidate trade-offs is in Appendix E.
5.1 Tale and Fine-tuning: How Does Pruning Interact with Training?
A natural question is whether pruning layers before or after fine-tuning harms the model’s ability to learn. Intuitively, one might expect that removing layers reduces representational capacity and thus limits downstream fine-tuning performance compared to baseline instruct-tuned models. Surprisingly, our experiments show the opposite: Tale not only preserves fine-tuning efficacy but in several cases improves both accuracy and efficiency.
We systematically explored four settings: (i) fine-tuning the base model (FT), (ii) applying Tale after fine-tuning (FT $\rightarrow$ Tale ), (iii) pruning first and then fine-tuning (Tale $\rightarrow$ FT), and (iv) pruning first then finetuning and then again pruning (BASE $\rightarrow$ Tale $\rightarrow$ FT $\rightarrow$ Tale ). Across reasoning and knowledge benchmarks, we consistently observed moderate but significant gains, especially on Winogrande (Table 2).
For example, pruning LLaMA-3.1 8B before fine-tuning reduced fine-tuning time by 2–2.5 GPU hours on an A100 (an 18.5% reduction) while simultaneously improving Winogrande performance by $2.4\%$. Iteratively applying pruning and fine-tuning allowed us to prune up to 8 layers achieving still higher accuracy (87.37%) than the full fine-tuned model (85.00%). Similarly, pruning the fully fine-tuned model yielded a 7-layer reduction while maintaining strong accuracy (86.66%).
On knowledge-heavy tasks such as MMLU and CommonsenseQA, the gains were smaller but consistent: pruned fine-tuned models retained or slightly exceeded the accuracy of their full fine-tuned counterparts, while requiring fewer parameters to optimize. This suggests that pruning can act as a regularizer, simplifying the optimization landscape by removing redundant layers.
Model Dataset Baseline Pruned Only FT Only Prune $\rightarrow$ FT FT $\rightarrow$ Prune (Prune $\rightarrow$ FT) $\rightarrow$ Prune
Perf. #D Perf. #D Perf. #D Perf. #D Perf. #D Perf. #D
Llama 3.1 8B Winogrande 53.83 0 56.67 4 85.00 0 87.06 4 86.74 7 87.37 8
MMLU 54.87 0 59.90 1 63.62 0 63.49 1 64.21 2 64.01 2
CommonQA 72.20 0 75.30 3 81.88 0 81.80 3 83.40 3 82.90 6
GSM8K 15.07 0 37.08 3 42.70 0 53.96 1 50.86 2 54.02 2
Qwen 0.5B Winogrande 49.86 0 51.88 5 50.43 0 50.43 5 50.49 2 52.49 9
MMLU 31.48 0 39.98 2 44.87 0 43.76 2 45.53 2 45.58 3
Overall, these results highlight an unexpected but consistent trend: pruning with Tale does not hinder fine-tuning but instead synergizes with it. Pruned models fine-tune faster, require fewer parameters to adapt, and are close to or better in performance than their full counterparts. Pruning finetuned models almost always improves performance. This shows that Tale pruning can be effectively interleaved with fine-tuning to create models that are both more accurate and computationally efficient.
6 Information theory and why pruned models perform better.
Our experiments show that selectively dropping layers improves accuracy for a targeted task. This seems counterintuitive: why could removing parts of a carefully trained model improve it?
Alemi et al. (2016); Tishby & Zaslavsky (2015) have used information theory (Shannon, 1948) to analyze how neural networks learn and represent data. We can use also this tool to explain the effects of our pruning algorithm. Fano & Hawkins (1961) define $\text{I}(\text{X};\text{Y})$, the mutual information between two random variables $X$ and $Y$, with the equation:
$\text{I}(\text{X};\text{Y})=\text{H}(\text{Y})-\text{H}(\text{Y}\mid\text{X})=\text{H}(\text{X})-\text{H}(\text{X}\mid\text{Y})=\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}p(x,y)\log\frac{p(x,y)}{p(x)\,p(y)}$ (1)
where $p(x,y)$ is the joint distribution of X and Y, and $p(x),p(y)$ are their marginals and where $\text{H}(\text{X})=-\sum_{x}p(x)\log p(x)$ is the Shannon (1948) entropy. $\text{I}(\text{X};\text{Y})$ measures how much knowing X reduces uncertainty about Y Tishby & Zaslavsky (2015); Shwartz-Ziv & Tishby (2017).
To apply these notions in our case, we estimate $\text{I}(\text{X}^{(l)},\text{Y})$ at each layer using trained probes as mutual information approximators. While this approach yields $\text{I}_{\text{probe}}(\text{X}^{(l)},\text{Y})$, an approximation of the true mutual information, probe-based estimates capture the linearly accessible task-relevant information at each layer, which is meaningful since many downstream tasks employ linear classification heads Belinkov (2022).
This estimate together with these three observations yields an information-theoretic explanation for why architectural pruning can improve model performance. (i) Certain layers in large pre-trained transformers decrease mutual information between the layer’s representation and the target task (Figure 3 and Table 7). (ii) From Tale we select the layer $\ell$ removed at the first iteration, and compare $\mathrm{I}\!\left(X^{(\ell+1)},Y\right)$ before deletion of layer $\ell$ to its value after deletion. Deleting this layer always increases the mutual information at the subsequent layer on the tasks, effectively preventing the previous layer, which acts as noise, from obstructing the flow of information through the network. Although our findings depend on the approximative nature of probe-based MI estimation, they provide evidence that certain layers in over-parameterized transformers act as information bottlenecks, which degrade rather than refine task-relevant representations. By removing layers that decrease mutual information with the target task, we enable representations with higher task-relevant information to flow directly to subsequent layers using the residual connection. This improves the information flow through the remaining architecture, results in representations that are more predictive of the target task, and yields improved accuracy at lower computational cost.
(a) ARC-Easy (Qwen 0.5B)
(b) BoolQ (Lucie 7B)
(c) BigBench (Llama 8B)
7 Discussion
We summarize five key observations below from our experiments.
1. Deleting later layers frequently improves performance on various tasks.
This challenges prior claims that early layers are largely redundant and more amenable to removal, instead highlighting their essential role in preserving core task-relevant representations. Even deleting many late layers does not reduce accuracy below baseline, whereas removing even a single early layer is often catastrophic (see Figure 4). All models exhibit similar behavior.
On the other hand, early layers appear crucial for task-solving, even though probing outputs at those layers does not yield interpretable responses. These results are potentially important for model interpretability. Plotting the performance degradation from ablating individual or consecutive layers, helps localize where specific task-solving abilities reside in the network.
2. Task dependence of layer importance.
Which layers improve or harm performance when removed is highly task dependent. Sometimes a single layer is critical: for instance, removing layer 25 of Llama-8B on CommonsenseQA causes a 50-point accuracy drop. Removing Llama’s layer 3 improves performance on GSM8K-hard but hurts MATH500; the revers happens when removing layer 11. Removing early layers (1–3) reduces accuracy to near zero on commonsense reasoning tasks (Figure 4), suggesting that certain early layers localize critical task-relevant information.
3. Structured task-specific patterns.
Although pruning is task-specific, related tasks often exhibit similar layer dependencies. Commonsense reasoning tasks (see Figure 4) show importance concentrated in comparable regions of the network. Mathematical reasoning tasks benefit from pruning one to three early layers (e.g., LLaMA layer 3, Mistral layers 6 and 22, Lucie layer 12), but not more (Figures 7, 8, 9). Commonsense and language tasks (ARC, BoolQ, CommonsenseQA, Winogrande, and BIG-Bench) benefit from deleting later layers (Tables 7, 9, 8). This suggests that later layers often play a decoding role for predictions into natural language, which reinforces point 1—pruning them doesn’t harm predictive capability.
We observe stronger pruning gains in reasoning-heavy tasks under zero-shot evaluation. All models showed notable accuracy boosts from one or two layer deletions on mathematical reasoning (e.g., Llama’s large gain on GSM8K-hard, and double-digit gains on MATH500 for Qwen, Mistral, and Lucie). By contrast, knowledge-intensive tasks exhibit more modest improvements (e.g., an 11% gain for Llama on BIG-Bench).
4. Model-specific pruning effects.
Different models display distinct pruning behavior. For example, pruned Lucie achieved a 101% gain on MMLU and double-digit gains on ARC-Challenge, CommonsenseQA, BoolQ and GSM8K-hard. While Qwen-7B, Llama-8B and Mistral share similar architecture and scale, they had modest gains on these data sets. Lucie also benefitted from more substantial pruning than the other models. Interestingly, Lucie was trained on a much smaller dataset (3T tokens vs. 15T for Llama and 13T for Qwen) and also tuned for French conversational proficiency Gouvert et al. (2025), suggesting that some layers for one of our tasks may encode multilingual-specific functionality.
This highlights intriguing interactions between pretraining and pruning efficiency. One hypothesis is that models trained close to their performance ceiling (via large-scale pretraining, instruction tuning or RLHF) yield smaller pruning gains, whereas models trained under limited objectives may benefit more. Notably, even the large corpus trained Qwen-0.5B showed strong pruning efficiency gains.
We experimented with producing pruned models for several tasks, and we saw that a single, general pruning improved speed up without much loss in accuracy across multiple tasks (Appendix E). This shows that judicious pruning can preserve generalization ability across several tasks.
5. Tale is versatile.
Tale can prune: base pre-trained models, instruction-tuned models (as we mainly do here), fine-tuned, and post-trained models with RLHF. Our results show that interleaving pruning and fine-tuning yields compounded benefits that may extend to other training regimes.
8 Conclusions
Tale is a generic toolkit for removing layers not needed for a certain task that yields increased performance and reduced computational costs. We have shown that it can profitably interact with further training or fine tuning, allowing users to further increase their models’ task specific performance. We have also shown that Tale can improve models both at small and larger scales. Tale can benefit high-throughput applications with time constraints–e.g. in multi-agent systems with task-specific agents, real-time recommendation engines, or interactive AI assistants. Tale can also help smaller companies or organizations that face critical trade-offs between model capability and computational efficiency use large language models at scale.
Appendix A Implementation Details
Hardware.
All experiments were conducted on 1 NVIDIA A100 GPU with 80GB memory.
Models.
Datasets for Tale pruning.
Pruning setup.
At each iteration, Tale evaluates all candidate single-layer deletions with respect to validation accuracy. The pruning threshold was defined as the baseline accuracy of the full model, ensuring that pruning never reduces performance relative to the original unpruned model. The iterative procedure terminates once no further layer removals satisfy this criterion.
Fine-tuning setup.
Evaluation.
We measured performance in terms of task accuracy on the test portions of datasets through automatic evaluation, with no human assessment involved, while inference efficiency was reported as relative speedup over the full baseline model. To evaluate MATH500 and GSM8K automatically without any human verification, we force the model to predict the final answer after the reasoning steps in a particular format and then calculate the accuracy. This could lower accuracy from what is expected; but since we used the same evaluation criteria for all the techniques and models and are interested in relative improvements or decreases in performance, the increase/decrease due to our methods from what might have been expected is justified.
Prompting.
Bigbench System Prompt ”You are a boolean expression evaluator. You must respond with exactly one word: either ’True’ or ’False’. Do not provide explanations, steps, or any other text. Only respond with ’True’ or ’False’.”
BOOLQ System Prompt ”You are a helpful assistant that answers True/False questions based on given passages. Read the passage carefully and determine if the question can be answered as True or False based on the information in the passage. ”Respond with only ’A’ for True or ’B’ for False.”
CommonQA System Prompt ”You are a helpful assistant that answers multiple-choice questions requiring commonsense knowledge and reasoning. Read each question carefully and select the most logical answer from the given options based on common knowledge and reasoning. Respond with only the letter of your chosen answer (A, B, C, D, or E).”
GSM8K System Prompt ”You are a math problem solver. Solve the given math problem step by step. ” ”Show your complete reasoning and calculations. ” ”At the end, write your final answer after ’####’ like this: #### [your final numerical answer]””
MMLU System Prompt ”You are a helpful assistant that answers multiple-choice questions across various academic subjects including humanities, social sciences, STEM, and professional fields. Read each question carefully and select the best answer from the given options. Respond with only the letter of your chosen answer (A, B, C, or D).”
Winogrande System Prompt You are a careful math problem solver. Show complete step-by-step reasoning and all calculations needed to arrive at the answer. Use clear, numbered or labeled steps so the reasoning is easy to follow. IMPORTANT (formatting): After the full reasoning, write the final answer on a new line by itself in exactly this format: ####
integer
Appendix B Number of parameters per layer for each model
Model LLaMA 3.1 8B Qwen 2.5 7B Mistral 7B Lucie 7B Qwen 2.5 0.5B
Parameters 218,112,000 233,057,792 218,112,000 192,946,176 14,912,384
Table 3: Model parameter counts comparison. LLaMA 3.1 8B, Mistral 7B and Lucie 7B has 32 layers, Qwen 2.5 7B has 28 layers and Qwen 2.5 0.5B has 24 layers.
Appendix C Intution behind Tale
(a) ARC-Challenge
(b) MMLU
Figure 5: Layer-wise output performance for LLaMA models: results when generating predictions from intermediate layers 1 through 32 on three different datasets.
Appendix D Results on few-shot learning on larger models
@@
Appendix E A tunable metric for finding accuracy vs. speed up optimization
$r_{A}=\frac{\text{Acc}(\text{Model})}{\text{Acc}(\text{Baseline})},\qquad\text{AE-HM}(\text{Model})=\frac{(1+\lambda^{2})r_{A}\,S}{\lambda^{2}S+r_{A}}=\frac{1+\lambda^{2}}{\frac{\lambda^{2}}{r_{A}}+\frac{1}{S}}$ (2)
$M_{\text{best-compromise}}=\arg\max_{i}\text{AE-HM}(M_{i})$ (3)
Figure 6: Relative Gain comparison across datasets. LLaMA $\beta=3$
Appendix F Deleted Layers in each Model and Benchmark
@@
Appendix G Common pruned layers model

### v2

@@ Appendix I LM-Eval as a sanity check
Appendix L Deleted Layers in each Model and Benchmark
@@
Dataset LLaMA 3.1 8B Qwen 2.5 7B
Baseline Best Model BSBA Baseline Best Model BSBA
Perf. Perf. #D Sp. Perf. #D Sp. Perf. Perf. #D Sp. Perf. #D Sp.
ARC-Easy 87.00 90.55 (+3.55%) 5 1.27 85.09 9 1.48 90.04 94.40 3 1.17 90.08 7 1.5
ARC-Challenge 75.86 78.62 (+2.76%) 4 1.26 76.90 7 1.41 86.55 92.00 (+5.45%) 2 1.3 86.55 6 1.37
BoolQ 85.00 86.20 (+1.20%) 3 1.10 83.00 8 1.36 81.9 83.9 4 1.3 80.00 7 1.45
MMLU 54.87 59.90 (+5.03%) 1 1.05 54.87 9 1.37 68.10 71.00 5 1.24 68.13 6 1.31
COMMONQA 72.20 75.30 (+3.10%) 3 1.21 70.10 8 1.34 80.30 84.4 2 1.12 80.50 6 1.27
WINOGRANDE 53.83 56.67 (+2.84%) 4 1.25 53.83 12 1.47 62.04 67.25 3 1.16 62.19 6 1.48
BIG-Bench 75.20 83.60 (+8.40%) 5 1.24 75.20 11 1.58 79.2 81.60 6 1.47 81.60 6 1.47
GSM8K-HARD 15.07 38.02 (+22.95%) 1 1.12 17.31 6 1.35 - - - - - - -
GSM8K* 42.15 56.56 (+14.41%) 1 48.14 3 - - - - - - -
Table 13: Comparison of LLaMA 3.1 8B and Qwen 2.5 7B across nine benchmarks. We report accuracy (%), number of layers dropped, and relative inference speedup. Best accuracy per dataset is in bold. (+X%) shows accuracy improvement from baseline.
Appendix M More on pruning and a common pruned layers model
Appendix N General Pruning results
@@
Appendix O TALE evaluation with perplexity
Model WikiText2 LAMBADA
Vanilla TALE Vanilla TALE
LLaMA 3.1 8B 24.6 24.9 28.1 28.9
Lucie 7B 46.1 36.4 52.5 43.8
Table 15: Perplexity scores for two models across WikiText2 and LAMBADA with Vanilla and TALE (sparisty 10%) configurations.
Model Method Sparsity WinoGr ARC-e ARC-c
Baseline 0% 41.2 51.7 40
LLaMA-2-7B SLEB 10% 18 (-56.3% $\downarrow$) 29 (-43.9% $\downarrow$) 28.8 (-28.0% $\downarrow$)
TALE 10% 56 (+35.9% $\uparrow$) 62.3 (+20.5% $\uparrow$) 50 (+25.0% $\uparrow$)
TALE 25% 51 (+23.8% $\uparrow$) 64.8 (+25.3% $\uparrow$) 47.6 (+19.0% $\uparrow$)
Baseline 0% 42 73.0 54.9
LLaMA-2-13B SLEB 10% 24.2 (-42.3% $\downarrow$) 43.5 (-40.4% $\downarrow$) 29.8 (-47.3% $\downarrow$)
TALE 10% 56.4 (+34.3% $\uparrow$) 77.3 (+5.9% $\uparrow$) 64.4 (+17.1% $\uparrow$)
TALE 25% 55.2 (+31.4% $\uparrow$) 75.3 (+3.2% $\uparrow$) 64.1 (+16.4% $\uparrow$)
Table 16: Accuracies (%) with Decoder Eval on zero-shot tasks for LLaMA-2-7B and LLaMA-2-13B
Appendix P Scores with relative increases in accuracy
Dataset LLaMA 3.1 8B (zero-shot) Qwen 2.5 7B (zero-shot)
Baseline Best Model BSBA Baseline Best Model BSBA
Perf. Perf. #D Sp. Perf. #D Sp. Perf. Perf. #D Sp. saved Perf. #D Sp.
ARC-Easy 87.00 90.55(+4.08% $\uparrow$) 5 -14.6% 87.82 8 -23.5% 91.01 91.82(+0.89% $\uparrow$) 2 -10.0% 90.91 5 -30.3%
ARC-Challenge 75.86 78.62(+3.63% $\uparrow$) 4 -11.7% 76.90 7 -20.5% 86.55 92.00(+6.45% $\uparrow$) 2 -6.7% 86.55 6 -19.9%
BoolQ 85.00 86.20(+1.40% $\uparrow$) 3 -8.8% 85.70 7 -17.6% 84.10 86.90(+3.22% $\uparrow$) 4 -13.3% 82.70 5 -23.2%
MMLU 54.87 59.90(+9.17% $\uparrow$) 1 -2.9% 54.87 9 -26.4% 68.10 71.00(+4.26% $\uparrow$) 5 -16.6% 68.13 6 -19.9%
CommonQA 72.20 75.30(+4.29% $\uparrow$) 3 -8.8% 73.10 6 -17.6% 80.30 84.40(+5.11% $\uparrow$) 2 -6.6% 80.50 6 -19.9%
Winogrande 53.83 56.67(+5.28% $\uparrow$) 4 -11.7% 53.83 12 -32.2% 62.04 67.25(+8.40% $\uparrow$) 3 -10.0% 62.19 6 -19.9%
BIG-Bench 75.20 83.60(+11.17% $\uparrow$) 5 -14.4% 75.20 11 -32.2% 79.20 81.60(+3.03% $\uparrow$) 6 -19.9% 81.60 6 -19.9%
GSM8K-HARD 15.07 37.08(+146.05% $\uparrow$) 1 -2.9% 35.0 4 -11.7% 7.9 27.0(+243.58% $\uparrow$) 2 -6.6% 19.1 4 -13.3%
Math500 20.50 26.00(+26.83% $\uparrow$) 1 -2.9% 26.00 3 -8.8% 18.00 27.00(+50.0% $\uparrow$) 2 -6.6% 21.00 4 -13.3%
Dataset Lucie 7B (zero-shot) Mistral 7B (zero-shot)
Baseline Best Model BSBA Baseline Best Model BSBA
Perf. Perf. #D Sp. Perf. #D Sp. Perf. Perf. #D Sp. Perf. #D Sp.
ARC-Easy 72.45 76.55(+5.66% $\uparrow$) 6 -18.1% 72.55 13 -39.2% 81.02 83.45(+4.23% $\uparrow$) 5 -15.4% 81.09 9 -27.7%
ARC-Challenge 48.00 53.79(+12.06% $\uparrow$) 7 -21.1% 51.38 11 -33.1% 72.20 74.83(+3.64% $\uparrow$) 6 -18.5% 72.41 8 -24.6%
BoolQ 53.70 77.50(+44.32% $\uparrow$) 5 -17.2% 60.60 19 -54.2% 80.36 83.20(+3.53% $\uparrow$) 6 -18.5% 80.60 10 -27.7%
MMLU 21.36 42.98(+101.2% $\uparrow$) 8 -24.1% 39.39 15 -45.2% 52.73 57.81(+9.63% $\uparrow$) 2 -6.2% 52.91 8 -24.6%
CommonQA 55.50 69.70(+25.59% $\uparrow$) 3 -9.1% 57.10 17 -48.2% 57.32 61.40(+7.12% $\uparrow$) 4 -12.3% 57.40 7 -21.5%
Winogrande 54.20 57.80(+6.64% $\uparrow$) 5 -27.1% 54.30 15 -45.2% 52.55 58.80(+11.53% $\uparrow$) 10 -30.7% 53.43 13 -40.0%
BIG-Bench 69.60 77.20(+9.84% $\uparrow$) 9 -27.1% 72.00 15 -45.1% 70.00 76.40(+9.14% $\uparrow$) 9 -28.0% 72.80 11 -33.8%
GSM8K-HARD 14.20 17.80(+25.35% $\uparrow$) 1 -3.1% 17.40 3 -9.1% 11.24 19.10(+69.92% $\uparrow$) 2 -6.2% 15.73 4 -12.3%
Math500 19.00 27.00(+42.11% $\uparrow$) 2 -6.0% 26.00 3 -9.1% 8.00 16.00(+100% $\uparrow$) 1 -3.1% 10.00 4 -12.3%
Dataset Qwen 2.5 0.5B (zero-shot)
Baseline Best Model BSBA
Perf. Perf. #D Sp. Perf. #D Sp.
ARC-Easy 40.00 60.91(+48.49% $\uparrow$) 3 1.1 48.36 5 1.4
ARC-Challenge 35.52 40.34(+13.57% $\uparrow$) 1 1.1 37.24 4 1.5
BoolQ 62.30 67.20(+7.87% $\uparrow$) 5 1.4 66.20 6 1.5
MMLU 31.48 39.97(+26.96% $\uparrow$) 2 1.1 33.90 5 1.4
CommonQA 42.40 49.10(+15.80% $\uparrow$) 2 1.3 44.00 3 1.4
Winogrande 49.86 51.88(+4.51% $\uparrow$) 5 1.3 49.87 17 3.9
BIG-Bench 72.40 73.60(+1.66% $\uparrow$) 2 1.2 73.60 2 1.2
GSM8K-HARD 6.74 11.24(+66.77% $\uparrow$) 1 1.2 8.99 2 1.2
Math500 8.00 12.00(+50.00% $\uparrow$) 1 1.1 9.00 2 1.1
Table 17: Performance comparison for LLaMA 3.1 8B, Qwen 2.5 7B, Lucie 7b, Mistral 7b and Qwen 2.5 0.5B under 0-shot evaluation on the 9 benchmarks with the training test split as in Table 2). We report accuracy (%), number of layers dropped, and relative inference speed in time.